1. solution of x2y5y-2x-2
2. solution of p0x-1(x-2x0)
3. solution of m5b2mb-2
4. solution of 4x2p0(3xp5-2x-2)
and how you arrive at the solutions thank you

Answer by stanbon(75887)

1. solution of x2y5y-2x^-2
= x^2y^6-2x^-2
= x^2y^6 -(2/x^2)
-------------------
2. solution of p^0x^-1(x-2x^0)
= (1/x)(x-2)
= 1-(2/x)
= (x-2)/x
= --------------------
3. solution of m^5*b^2mb^-2
= m^5b^0
= m^5
----------------
4. solution of 4x^2p^0(3xp^5-2x^-2)
= 4x^2(3xp^5-(2/x^2))
= 12x^3p^5 - 8
=====================